OpenStudy (anonymous):

You want to build a right triangular garden on the side of your house. The hypotenuse of the triangle measures 8 feet. The other two sides are equal. What are the lengths of the two sides?

OpenStudy (anonymous):

let the other side = x so x^2 + x^2 = 8^2 2x^2 = 64 x^2 = 32 x = 4√2

OpenStudy (anonymous):

oh ok thnx

OpenStudy (anonymous):

ya no problem @iceywheniplease7